The management of Consumers Mfg. would like to purchase a specialized production machine for $222,000. The machine is expected to last three years, with a salvage value of $5,000. Annual maintenance costs will total $30,000, and annual labor and material savings are predicted to be $140,000. The company's required rate of return in 15%. Find the NPV of this investment.

1. $32,442

2. $6,404

3. $11,564

4. $28,210
Answer #1
Net annual cash inflow = Annual labor and materials saving - Annual maintenance cost = 140000 - 30000 110000
Present value of net annual cash inflow [ Net annual cash inflow * Present Value Annuity Factor @ 15% for 3 years = 110000 * 2.283225 ]    251154.75
(+) Present value of salvage value [ Salvage value * Present value interest factor @ 15% of 3rd year = 5000 * 0.657516 ] 3287.58
(-) Initial cost of machine 222000
Net present value (NPV) 32442.33
Rounded off 32442 Option 1